# The global communication pathways of the human brain transcend the cortical - subcortical - cerebellar division

### v1

@@ Methods
fMRI data acquisition
@@
In this work we employed the connectivity dataset from Ref. [Ji_Mapping_2019, ], which based on the Human Connectome Project (HCP) with 32 participants (“MGH HCP Adult Diffusion,” 16 females, 16 males) collected at the Massachusetts General Hospital van2013wu .
Data Acquisition.
The whole-brain echo-planar imaging was conducted with a modified 3T Siemens Skyra (Connectome Skyra) system with a 32 channel head coil applying a time to repetition (TR)=720ms, time to echo (TE)=33.1ms, flip angle=52, bandwidth=2,290 Hz/pixel, in-plane field of view (FOV)=$208\times 180$mm, 72 slices, and 2.0mm isotropic voxels, with a multi-band acceleration factor of 8, Ref. [ugurbil2013pushing, ]. For more details see Ref. [Ji_Mapping_2019, ].
Functional networks and subcortical / cerebellar ROI assignment.
The functional network calculation and the subcortical / cerebellar ROI assignment was previously performed by Ref. [Ji_Mapping_2019, ]. In detail, subject-specific functional connectivity matrices were calculated by correlating the individual cortical BOLD signals with Pearson correlation. By averaging across all subject matrices, a group-averaged functional connectivity (FC) matrix was constructed. Cortical resting-state functional networks were obtained by applying the Louvain algorithm to this group-averaged functional connectivity matrix Blondel_FastUnfolding_2008 . Next, subject-wise FC matrices indicating the correlation between the 360 cortical ROIs and 31870 subcortical / cerebellar grayordinates spanning over the entire CIFTI space were calculated and averaged to obtain a group-representative cortical–subcortical / cerebellar FC matrix. These grayordinates were assigned to cortical functional networks with the highest correlation and merged together based on this partition. At the same time, the parcels were constrained to major subcortical structures as determined by Freesurfer, thus adhering to the general subcortical anatomy. This resulted in a partition of 233 subcortical and 125 cerebellar ROIs based on the detected cortical functional networks.
Whole-brain structural network construction
@@
The structural connectomes of the participants were averaged to generate a representative connectome. For the DWI a high-quality protocol was applied (i.e. b-value of 10,000 s/mm2, high-angular resolution, and a multi-slice approach). The data are publicly included within the Lead-DBS software package and were already preprocessed setsompop2013pushing ; horn2017connectivity . A generalised q-sampling imaging algorithm (DSI Studio; http://dsi-studio.labsolver.org) was used for the data processing. SPM 12 was used to segment and co-register the data. 200 000 fibres were examined for each participant using Gibbs’ tracking approach kreher2008gibbs and being restricted by a co-registered white-matter mask. The fibres were standardised into MNI space via DARTEL transforms ashburner2007fast ; horn2016toward . The cortex was parcelled into 360 regions (180 per hemisphere) according to Ref. [glasser2016multi, ].
@@
The cortex and cerebellum are contra-laterally connected via the cortico-ponto-cerebellar (CPC) and the cerebello-thalamo-cortical (CTC) pathways Ramnani_CorticoCereb_2006 ; Palesi_Pathways_2015 ; Palesi_Contralateral_2017 ; DAngelo_Physiology_2018 . To only display direct anatomical connections, initial cortical-cerebellar connections were filtered as diffusion imaging cannot distinguish between direct and indirect pathways. This resulted in the filtering of 8.3% of initial connections. Moreover, ipsilateral connections between the cerebellum and the brainstem (pons) / thalamus (total of 2.0% of initial connections) and direct inter-hemispheric cerebellar connections (as the cerebellar hemispheres are indirectly connected via the vermis DAngelo_Physiology_2018 ; 2.6% of initial connections) were filtered.
@@
The individual SC matrices for the 32 participants were averaged into one population brain-wide structural connectivity network, represented as the weighted global connectivity matrix $W$ of $718\times 718$ (ROIs). To perform a graph analysis, commonly $W$ is binarised by applying a hard threshold $\theta$ such that only the connections with a weight larger than $\theta$ are conserved. That is, the binary adjacency matrix $A$ is defined as $A_{ij}=1$ if $W_{ij}>\theta$, and $A_{ij}=0$ if $W_{ij}<\theta$. The value of $\theta$ then controls the final number (density) of connections. This typical hard thresholding leads to a bias. Since one of the limitations of tractography is that it overestimates shorter connections, the hard thresholding favours short-range links in detriment of long-range ones roberts2017consistency ; betzel2019distance . Nevertheless, the existence of long-range connections has been shown in tract-tracing studies ercsey2013predictive ; horvat2016spatial and are thought to serve important functional roles betzel2019distance as these connections cause a higher metabolic cost for the organism bullmore2012economy . To threshold the weighted matrix $W$ while also maintaining long-range connections, we introduce the “relative weight distance-dependent” (RWDD) thresholding. The goal of RWDD is to conserve the original distribution of connection lengths by performing an adaptive threshold such that $\theta=\theta_{d}$ depends on the distance $d(i,j)$ between two ROIs. Here, $d(i,j)$ will represent the Euclidean distance between the centers of mass of two ROIs $i$ and $j$, but it could be replaced by the fiber-length of the tracts when this information is available.
All links of the initial SC matrix were binned into N bins according to their length. This information is taken from a distance matrix, representing the distances between all nodes. Thus, each bin contains similarly long connections. Next, the relative weight of each bin is calculated, i.e., the number of connections in the bin compared to all connections in the initial SC matrix. Finally, the bins are reduced such that this relative weight is maintained and the number of final links for the thresholded SC matrix is reached. While doing so, only the strongest connections of the bins are retained.
Consequently, the RWDD thresholding was applied with a target density of 0.2. While targeting for a low density to reduce false-positive connections, unconnected components (i.e., ROIs with no connections to other ROIs) could be created. To account for this, strong connections of otherwise unconnected ROIs were being kept. If an ROI had the risk of becoming disconnected, links with a weight of 1.8 standard deviations above the mean were kept. This resulted in 12 additional links conserved which would initially violate the RWDD thresholding criteria.
@@
Several network measures were applied to characterise the brain-wide structural connectome here investigated.
@@
The density $\rho(G)$ of a network $G$, is the fraction between the number of links $L$ present and the maximum possible number of links $L_{max}=\frac{1}{2}N(N-1)$, where $N$ is the number of ROIs in a network. Similarly, the connection densities between modules in a network (e.g., cortex and subcortex) were calculated as the fraction of links $L_{\alpha\beta}$ present between modules $\alpha$ and $\beta$, and the maximum possible number of links between them $N_{\alpha}\,N_{\beta}$ where $N_{\alpha}$ and $N_{\beta}$ are the number of nodes in each module.
The degree of a node $k_{i}$ is the number of nodes to which node $i$ connects.
@@
The participation coefficient $p_{i}$ is a measure how a node distributes its links among modules Guimera_RolesJSM_2005 . Here we applied the normalised participation index defined in Ref. [Klimm_Roles_2014, ], as it takes into account the relative sizes of communities. The participation of node $i$, in a network divided into $M$ modules is defined as:
$p_{i}=1-\frac{M}{\sqrt{M-1}}\sigma\left(P_{i}\right)\,,$ (1)
@@
where $P_{i}$ is the participation vector of node indicating the probability that node $i$ belongs to each of the $M$ modules, and $\sigma\left(P_{i}\right)$ is the standard deviation of the participation vector. When $\sigma\left(P_{i}\right)=0$ indicates that the node has only links within the own community, whereas a value of $\sigma\left(P_{i}\right)=1$ indicates the node is equally likely connected to all the modules in the network.
@@
A Rich-club consists of supramodule of highly interconnected hubs Zhou_RichClub_2004 . The quantitative definition of a rich-club is an undisclosed problem since, in most cases, no unique set of nodes can be strictly defined as ‘the rich-club’. Instead, the presence or not of a rich-club is discriminated evaluating the $k$-density $\Phi(k^{\prime})$ function, which is the internal link density between the nodes with $k$ larger than a given $k^{\prime}$ such that:
$\Phi(k^{\prime})=\frac{L_{k^{\prime}}}{N_{k^{\prime}}(N_{k^{\prime}}-1}\,,$ (2)
@@
where $N_{k^{\prime}}$ is the number of nodes with degree $k>k^{\prime}$ and $L_{k^{\prime}}$ is the number of links between them. Iteratively computing $\Phi(k^{\prime})$ for $k^{\prime}=0,1,2,\ldots,k_{max}$, if the function monotonically increases, then we can assure a rich-club is present in the network. To perform a comprehensive analysis, instead of taking a single threshold, we studied the resulting rich-club organization for different thresholds in the range of $\Phi(k^{\prime})=[0.8,1.0]$. For comparative purposes, $k$-density was also estimated out of an ensemble of 100 random graphs conserving the degree distribution, using the link-switching method Katz_Rewiring_1957 .
@@
Community detection. We applied the Leiden community detection algorithm Traag_Leidenalg_2019 to the brain-wide structural matrix. To account for different resolution parameters, we first applied the ‘RB’ null-model Reichardt_Partitioning_2007 . For each step of size of 0.05 in the range of $\gamma\in[0.6,1.4]$, the algorithm was run 100 times and the partition with the highest quality value selected. Finally, the partition of this set with the highest modularity index Newman_Modularity_2004 was chosen, see Supplementary Fig. S1.
@@
Integration and segregation of the brain-wide network were estimated using the definitions introduced in Zamora_Hubs_2010 . This approach accounts for the integration capacity of a group of nodes and for the segregation (loss of communication) suffered by a network, after selected nodes have been lesioned. These are defined in general and can be evaluated under different dynamical or propagation models. For the exploratory purposes of the present study, we employed an analytical account of the dynamical influence that one node $j$ exerts over another $i$, by measuring the newtork response measured at $i$ to an stimulus applied at $j$.
Stimulus-response matrices
. Assuming the dynamics of the network (represented by its connectivity matrix $A$) are governed by a continuous leaky-cascade (a multivariate autoregressive process), the temporal evolution of the system is given by:
$\dot{\boldsymbol{x}}=-\frac{\boldsymbol{x}}{\tau}+A\,\boldsymbol{x}\,,$ (3)
where $x(t)=[x_{1}(t),x_{2}(t),x_{3}(t),\ldots,x_{N}(t)]^{T}$ is the (column) vector of the temporal states $x_{i}(t)$ for nodes $i=1,2,\ldots,N$, and $\tau$ is a leakage or a dissipation time-constant. In this linear system, the temporal response $R_{ij}(t)$ of node $i$ at times $t>0$ to a unit stimulus applied at node $j$, at time $t=0$, can be analytically estimated Gilson_DynCom_2019 ; Zamora_Chaos_2024 . Assuming that initially all nodes receive a unit stimulus, the temporal response matrix is obtained as:
@@
where $J_{ij}=-\frac{\delta_{ij}}{\tau}+A_{ij}$ is the Jacobian matrix of the linear dynamical system and $J^{0}_{ij}=-\frac{\delta_{ij}}{\tau}$ corresponds to the trivial leakage through a node, due to a perturbation applied on itself. This pair-wise response encompasses all network effects from $j$ to $i$ acting at different time scales along all recurrent paths of different lengths. If $\lambda_{max}$ is the largest eigenvalue of the connectivity matrix $A$, the responses will converge to zero after a transient response only if $\tau<1/\lambda_{max}$. Otherwise, if $\tau>1/\lambda_{max}$, the system diverges and so do the responses measured. For the calculations here, we considered $\tau=0.5\,(1/\lambda_{max})$.
So far, $R_{ij}(t)$ represents the temporal evolution of the responses. For the estimation of integration and segregation, the total response of node $j$ to $i$ were obtained as the accumulated response over time, by computing its integral over time (or area-under-the-curve from $t=0$ to $t=\infty$),
@@
In the following, we will refer this $N\times N$ matrix $R$ as the ‘pair-wise response matrix’.
@@
Here, we define integration not as a global property of a network, but as the capacity of one node, or of a set of selected nodes, to integrate the information in the network Zamora_Hubs_2010 . Let $X$ be the set of all nodes in a network and $H\subset X$ a subset of arbitrarily chosen nodes. Then, the integration capacity $I(H)$ is defined as the joint response by the nodes in subset $H$, conditional to the simultaneous stimuli applied at all other nodes, the complementary set $X-H$. Formally,
@@
where $\mathcal{R}$ is a generic response function. Note that the nodes in $H$ are not stimulated, instead, they “passively listen” to the stimuli applied everywhere except on them. This definition could be applied for an arbitrary dynamical processes running on the network. In here, we consider the linear dynamical process above. Given the vector of unit stimuli $\mathbf{s}(X-H)$ where $s_{i}=0$ if $i\in H$ and $s_{i}=1$ for $i\in X-H$, the corresponding conditional temporal responses $R(t,X-H)$ are
$R(t;X-H)=\left(e^{Jt}-e^{J^{0}t}\right)\,\mathbf{s}(X-H)$ (7)
@@
$I(H)=\sum_{i\in X-H;\,j\in H}R_{ij}(X-H).$ (9)
For the results in Fig. 5b, the $H$ is considered as the set of 30 ROIs with largest degree $k$, and $X-H$ are the remaining 688 ROIs. The hub sets $H$ are always ranked based on their global degree, as in Fig. 4a, but the selection for the different cases are based on the community the ROIs belong to, in either the anatomical or the network-based division. Hence, the colored bars all correspond to the integration capacity of 30 hubs, but only for those that belong to a specific module. The results in Fig. 5c are computed equivalently but considering hub sets $H$ of increasing size, from $|H|=1$ to $|H|=170$, adding one hub at a time to the set $H$.
@@
Here, we define segregation as the loss of communication suffered between the modules of a network, after a node or a set of nodes has been lesioned Zamora_Hubs_2010 . Let $\mathcal{P}$ be a partition of a network into $M$ modules, where $C_{m}$ with $m=1,2,\ldots,M$ being the corresponding sets of nodes. We define modular integration $I_{\mathcal{P}}(X)$ of the system $X$, as the amount of communication between modules. In this case, we evaluate $I_{\mathcal{P}}(X)$ as the sum of pairwise responses of the nodes in one module $n$, to the stimuli applied at the nodes of module $m$. Let $R_{nm}$ be the reduced modular response matrix of size $M\times M$, evaluated summing the responses of all nodes in $n$, to stimuli applied at $m$,
$R_{nm}(X)=\sum_{i\in C_{n};\,j\in C_{m};\,n\neq m}R_{ij}\,.$ (10)
Then, modular integration is calculated as the sum of the extradiagonal entries of matrix $R_{nm}$, such that:
$I_{\mathcal{P}}(X)=\sum_{n,m=1;\,n\neq m}^{M}R_{nm}\,.$ (11)
Assume now that a set of nodes $H\subset X$ is lesioned, i.e., deleted from the network, and modular integration is recomputed for the same partition but in the lesioned network, namely, $I_{\mathcal{P}}(X-H)$. Segregation is thus defined as the (normalised) loss of modular integration in the network after the lesion of set $H$, compared to the modular integration in the original ‘healthy’ network:
@@
For the segregation results in Fig. 5, first the reduced modular response matrix $R_{nm}(X)$ and the corresponding modular integration $I_{\mathcal{P}}(X)$ for the original–unlesioned–network were evaluated in both the anatomical and the network divisions. Then, targeted lesions were applied for various cases and the resulting modular integration $I_{\mathcal{P}}(X-H)$ was recomputed. This implied to apply a lesion to the network for the chosen set of nodes $H$ by deleting the corresponding rows and columns in the connectivity matrix $A$, and to recompute the resulting pair-wise response matrix for the lesioned connectivity, with stimulus vector $\mathbf{s}(X-H)$ where $s_{i}=0$ if $i\in H$ and $s_{i}=1$ for $i\in X-H$. Figure 5e reports the losses in modular integration between the modules, $R_{nm}(X-H)\,/\,R_{nm}(X)$, displayed as percentages, after the 30 nodes with largest degree had been lesioned. Figure 5f reports the segregation values when lesions were performed targeting either the 30 global hubs (black), or the 30 nodes with largest degree in each community (colored bars). Finally, the results were computed equivalently, but applying lesions of different sizes, from $|H|=1$ to $|H|=170$, by orderly including the nodes with largest degree, one-by-one, to the lesioned set $H$.

### v2

@@ Methods
Whole-brain structural network construction
To create the whole-brain structural connectome we employed diffusion imaging and tractography that was already preprocessed and publicly available via the Lead-DBS software package Horn and Kühn (2015). This tractography data is based on high-quality diffusion imaging from the Human Connectome Project Van Essen et al. (2013); Setsompop et al. (2013). We then parcellated the data following Ji et al. (2019), which includes subcortical structures and the cerebellum.
@@
Data of 32 participants from the Human Connectome Project (HCP) with (“MGH HCP Adult Diffusion,” 16 females, 16 males) was collected at the Massachusetts General Hospital Van Essen et al. (2013).
@@
For the diffusion weighted imaging, a high-quality protocol Setsompop et al. (2013) was applied (i.e. b-value of 10,000 $s/mm^{2}$, high-angular resolution, and a multi-slice approach). A generalized q-sampling imaging algorithm (DSI Studio; http://dsi-studio.labsolver.org) was used for the data processing. SPM 12 was used to segment and co-register the data. 200,000 fibres were examined for each participant using Gibbs’ tracking approach Kreher et al. (2008) and being restricted by a co-registered white-matter mask. The fibres were standardized into MNI space via DARTEL transforms Ashburner (2007); Horn and Blankenburg (2016).
Cortical, subcortical and cerebellar ROI assignment.
We employed the brain-wide parcellation developed by Ji et al. (2019). The cortex was first parcellated into 360 regions (180 per hemisphere) according to Glasser et al. (2016). The subcortical and cerebellar ROI assignment was performed based on clustering of resting-state functional connectivity, see Ji et al. (2019) for details. In summary, subject-specific functional connectivity (FC) matrices were calculated by correlating the individual cortical BOLD signals with Pearson correlation. By averaging across all subject matrices, a group-averaged FC matrix was constructed. Cortical resting-state functional networks were obtained by applying the Louvain algorithm to this group-averaged FC matrix Blondel et al. (2008). Next, subject-wise FC matrices indicating the correlation between the 360 cortical ROIs and 31,870 subcortical and cerebellar grayordinates spanning over the entire CIFTI space were calculated and averaged to obtain a group-representative cortical – subcortical / cerebellar FC matrix. These grayordinates were assigned to cortical functional networks with the highest correlation and merged together based on this partition. At the same time, the parcels were constrained to major subcortical structures as determined by Freesurfer, thus adhering to the general subcortical anatomy. This resulted in a partition of 233 subcortical and 125 cerebellar ROIs based on the detected cortical functional networks.
@@
The cortex and cerebellum are contra-laterally connected via the cortico-ponto-cerebellar and the cerebello-thalamo-cortical pathways Ramnani (2006); Palesi et al. (2015, 2017); D’Angelo (2018). To only display direct anatomical connections, ($i$) cortico-cerebellar tracts were filtered as diffusion imaging cannot distinguish between direct and indirect pathways. This resulted in the filtering of 8.3% of initial connections. ($ii$) Ipsilateral tracts between the cerebellum and the brainstem (pons), ($iii$) ipsilateral tracts between cerebellum and thalamus (total of 2.0% of initial connections) where filtered out, as well as ($iv$) direct inter-hemispheric cerebellar tracts (2.6% of initial tracts), as the cerebellar hemispheres are indirectly connected via the vermis D’Angelo (2018).
@@
The individual SC matrices for the 32 participants were averaged into one population brain-wide structural connectivity network, represented as the weighted global connectivity matrix $W$ of $718\times 718$ (ROIs). To perform graph analysis, usually $W$ is binarized by applying a hard threshold $\theta$ such that only connections with a weight larger than $\theta$ are conserved. A binary adjacency matrix $A$ is thus defined as $A_{ij}=1$ if $W_{ij}>\theta$, and $A_{ij}=0$ if $W_{ij}<\theta$. The value of $\theta$ controls for the final number (density) of connections. This typical hard thresholding leads to a bias since tractography overestimates shorter connections. Thus, the hard thresholding favours short-range links in detriment of long-range ones Roberts et al. (2017); Betzel et al. (2019).
To obtain a binary adjacency matrix $A$ preserving long-range connections, we introduce the “relative weight distance-dependent” (RWDD) thresholding. The goal of RWDD is to conserve the original distribution of connection lengths by performing an adaptive threshold such that $\theta=\theta_{d}$ depends on the distance $d(i,j)$ between two ROIs. Here, $d(i,j)$ will represent the Euclidean distance between the centers of mass of two ROIs $i$ and $j$, but it could be replaced by the fiber-length of the tracts when this information is available. First, all links (non-zero values of the population average matrix $W_{ij}$) are sorted into $n$ bins according to the distance between ROIs. Thus, each bin contains all links of similar length $d$. The fraction of links falling in a bin (to the total number of links) is calculated. Finally, weaker links are filtered out to meet the desired target number of links in the network by applying an adaptive threshold to each bin individually, such that the fractions of links of lengths $d$ are conserved.
The RWDD thresholding was applied with a target density for $A$ of 0.2. While targeting for a low density reduces the number of false-positive connections, unconnected components could appear (i.e., ROIs with no connections to other ROIs). To correct for this, strong connections of otherwise unconnected ROIs were kept. If an ROI had the risk of becoming disconnected, links with a weight of 1.8 standard deviations above the mean were kept. This resulted in 12 additional links conserved which would initially violate the adaptive RWDD thresholding criteria.
@@
Given the adjacency matrix $A$ for a network of $N$ nodes and $L$ links, the density $\rho=L\,/\,L^{0}$ of the network is the fraction between the number of links $L$ present and the maximum possible number of links $L^{0}=\frac{1}{2}N(N-1)$. If the nodes of a network are partitioned into M groups or modules, e.g., the division of the ROIs into cortical, subcortical and cerebellar, each of $N_{\alpha}$ nodes and $L_{\alpha}$ links, the internal connection of each group is $\rho_{\alpha}=L_{\alpha}\,/\,L^{0}_{\alpha}$ where $L^{0}_{\alpha}=\frac{1}{2}N_{\alpha}(N_{\alpha}-1)$. The connection density (probability) between two groups (e.g., cortical and subcortical ROIs) is $\rho_{\alpha\beta}=L_{\alpha\beta}\,/\,L_{\alpha\beta}^{0}$, with $L_{\alpha\beta}^{0}=N_{\alpha}\,N_{\beta}$ being the number of all possible connections between groups $\alpha$ and $\beta$.
@@
The degree $k_{i}$ of a node is the number of nodes to which node $i$ connects. The participation coefficient $p_{i}$ is a measure of how a node distributes its links among modules Guimerà and Amaral (2005). Here we employed the unbiased participation index defined in Klimm et al. (2014), which takes into account the relative sizes of communities. In a network divided into $M$ modules, the participation vector $\mathbf{p}_{i}$ of the network is defined as the vector of length $M$ where $P_{i,m}$ represents the probability of a node $i$ to connect to module $m$. Then, the participation index of $i$ is:
$p_{i}=1-\frac{M}{\sqrt{M-1}}\sigma\left(\mathbf{p}_{i}\right)\,,$ (1)
@@
where $\sigma\left(P_{i}\right)$ is the standard deviation of the participation vector. When $p_{i}=0$, it indicates that $i$ is only linked to other nodes within the community it belongs, whereas a value of $p_{i}=1$ indicates that $i$ is equally likely connected across all modules in the network.
@@
A rich-club consists of supra-module of highly interconnected hubs Zhou and Mondragón (2004). The quantitative definition of a rich-club is an undisclosed problem since (in most cases) no unique set of nodes can be strictly defined be the rich-club of a network. However, the presence or not of a rich-club can be discriminated evaluating the $k$-density $\Phi(k^{\prime})$ function, which is the internal link density between the nodes with $k$ larger than a given $k^{\prime}$ such that:
$\Phi(k^{\prime})=\frac{L_{k^{\prime}}}{N_{k^{\prime}}(N_{k^{\prime}}-1)}\,,$ (2)
@@
where $N_{k^{\prime}}$ is the number of nodes with degree $k>k^{\prime}$ and $L_{k^{\prime}}$ is the number of links between them. Iteratively computing $\Phi(k^{\prime})$ for $k^{\prime}=0,1,2,\ldots,k_{max}$, if the function monotonically increases, then we can assure a rich-club is present in the network. To perform a comprehensive analysis, instead of taking a single threshold, we studied the resulting rich-club organization over various thresholds in the range of $\Phi(k^{\prime})=[0.8,1.0]$. For comparative purposes, $k$-density was also estimated out of an ensemble of 100 random graphs conserving the degree distribution, using the link-switching method Katz and Powell (1957).
@@
Community detection. We applied the Leiden community detection algorithm Traag et al. (2019) to the brain-wide structural matrix. To account for different resolution parameters, we first applied the ‘RB’ null-model Reichardt and Bornholdt (2007). For each step of size of 0.05 in the range of $\gamma\in[0.6,1.4]$, the algorithm was run 100 times and the partition with the highest quality value selected. Finally, the partition of this set with the highest modularity index Newman and Girvan (2004) was chosen, see Supplementary Fig. S1.
Community detection was performed with the leidenalg Python package, https://pypi.org/project/leidenalg/. All other network analyses were carried using the pyGAlib Python package, https://github.com/gorkazl/pyGAlib.
@@
Integration and segregation of the brain-wide network were estimated using the definitions introduced in Zamora-López et al. (2010). This approach accounts for the propagation of external stimuli throughout the network and the subsequent collective responses by groups of brain regions. While such propagation and collective responses could be broadly represented, for the exploratory purposes of the present study, we employed a linear dynamical approach that allows to analytically estimate the influence that one node $j$ exerts over another $i$ encompassing all possible paths, of all lengths, from $j$ to $i$.
Stimulus-response matrices.
Assuming the dynamics of the network (represented by its connectivity matrix $A$) are governed by a linear multivariate autoregressive process, the temporal evolution of the system is given by:
$\dot{\bm{x}}=-\frac{\bm{x}}{\tau}+A\,\bm{x}\,,$ (3)
where $x(t)=[x_{1}(t),x_{2}(t),x_{3}(t),\ldots,x_{N}(t)]^{T}$ is the (column) vector of the temporal states $x_{i}(t)$ for all nodes, and $\tau$ is a leakage or dissipation time-constant. In this linear system, the temporal response $R_{ij}(t)$ of node $i$ at times $t>0$ to a unit stimulus applied at node $j$, at time $t=0$, can be analytically estimated Gilson et al. (2019); Zamora-López and Gilson (2024). Assuming that initially all nodes receive a unit stimulus, the temporal pair-wise responses are obtained as:
@@
where $J_{ij}=-\frac{\delta_{ij}}{\tau}+A_{ij}$ is the Jacobian matrix of the linear dynamical system and $J^{0}_{ij}=-\frac{\delta_{ij}}{\tau}$ corresponds to the trivial leakage through a node, due to a perturbation applied on itself. This pair-wise response encompasses all network effects from $j$ to $i$ acting at different time scales along all recurrent paths of different lengths. If $\lambda_{max}$ is the largest eigenvalue of the connectivity matrix $A$, the responses will converge to zero after a transient response only if $\tau<1/\lambda_{max}$. Otherwise, if $\tau>1/\lambda_{max}$, the system diverges. For the calculations here, we considered $\tau=0.5\,(1/\lambda_{max})$.
So far, $R_{ij}(t)$ represents the temporal evolution of the pair-wise responses. For the estimation of integration and segregation, the total response of node $j$ to $i$ were obtained as the accumulated response over time, by computing its integral over time (or area-under-the-curve) from $t=0$ to $t=\infty$),
@@
In the following, we will refer to this $N\times N$ matrix $R$ as the ‘pair-wise response matrix’. The stimulus-response dynamics and the matrices $R$ were obtained using the SiReNetA Python package, https://github.com/mb-BCA/SiReNetA/tree/master.
@@
Here, we define integration as the sensitivity of one node (or of a set of selected nodes) to external stimuli applied to the network Zamora-López et al. (2010). Let $X$ be the set of all nodes in a network and $H\subset X$ a subset of arbitrarily chosen nodes. Then, the integration capacity $I(H)$ is defined as the joint response by the nodes in subset $H$, conditional to the simultaneous stimuli applied at all other nodes, the complementary set $X-H$. Formally,
@@
where $\mathcal{R}$ is a generic response function. Note that the nodes in $H$ are not stimulated, instead, they “passively listen” to the stimuli applied everywhere except on them. This definition could be applied for an arbitrary dynamical processes running on the network. Here, we consider the linear dynamical process in Eq. (3). Given the vector of unit stimuli $\mathbf{s}(X-H)$ with $s_{i}=0$ if $i\in H$, and $s_{i}=1$ for $i\in X-H$, the temporal evolution of the response matrix is
$R(t;X-H)=\left(e^{Jt}-e^{J^{0}t}\right)\,\mathbf{s}(X-H)\,,$ (7)
@@
$I(H)=\sum_{j\in H\atop i\in X-H}R_{ij}(X-H).$ (9)
For the results in Fig. 5b, $H$ is considered as the set of 30 ROIs with largest degree $k$, and $X-H$ are the remaining 688 ROIs. The hub sets $H$ are always ranked based on their global degree, as in Fig. 4a, but the selection for the different cases are based on the community the ROIs belong to, in either the anatomical or the network-based division. Hence, the colored bars all correspond to the integration capacity of 30 hubs, but only for those that belong to a specific module. The results in Fig. 5c are computed equivalently but considering hub sets $H$ of increasing size, from $|H|=1$ to $|H|=170$, adding one hub at a time to the set $H$.
@@
We define segregation as the loss of communication (influence) suffered between the modules of a network, after a node or a set of nodes has been lesioned Zamora-López et al. (2010). Let $\mathcal{P}$ be a partition of a network into $M$ modules and $C_{m}$ the set of nodes in module $m$. We define modular integration $I_{\mathcal{P}}(X)$ of the system $X$, as the amount of communication between modules. In this case, we evaluate $I_{\mathcal{P}}(X)$ as the sum of pairwise responses of the nodes in one module $n$, to the stimuli applied at the nodes of module $m$. Let $R_{nm}$ be the reduced modular response matrix of size $M\times M$, evaluated summing the responses of all nodes in $n$, to stimuli applied at $m$,
$R_{nm}(X)=\sum_{i\in C_{n}}\sum_{j\in C_{m}\atop m\neq n}R_{ij}\,.$ (10)
Then, modular integration is calculated as the sum of the extra-diagonal entries of matrix $R_{nm}$, such that:
$I_{\mathcal{P}}(X)=\sum_{n=1}^{M}\sum_{m=1\atop m\neq n}^{M}R_{nm}\,.$ (11)
Assume now that a set of nodes $H\subset X$ is lesioned, i.e., deleted from the network, and modular integration is recomputed for the same partition but in the lesioned network, namely, $I_{\mathcal{P}}(X-H)$. Segregation is thus defined as the (normalized) loss of modular integration in the network after the lesion of set $H$, compared to the modular integration in the original ‘healthy’ network:
@@
For the segregation results in Fig. 5, first the reduced modular response matrix $R_{nm}(X)$ and the corresponding modular integration $I_{\mathcal{P}}(X)$ for the original–unlesioned–network were evaluated in both the anatomical and the network divisions. Then, targeted lesions were applied for various cases and the resulting modular integration $I_{\mathcal{P}}(X-H)$ was recomputed. This implied to apply a lesion to the network for the chosen set of nodes $H$ by deleting the corresponding rows and columns in the connectivity matrix $A$, and to recompute the resulting pair-wise response matrix for the lesioned connectivity, with stimulus vector $\mathbf{s}(X-H)$ where $s_{i}=0$ if $i\in H$ and $s_{i}=1$ for $i\in X-H$. Figure 5e reports the losses in modular integration between the modules, $R_{nm}(X-H)\,/\,R_{nm}(X)$, displayed as percentages, after the 30 nodes with largest degree had been lesioned. Figure 5f reports the segregation values when lesions were performed targeting either the 30 global hubs (black), or the 30 nodes with largest degree in each community (colored bars). Finally, for the results in Fig. 5g were computed equivalently, but applying lesions of different sizes, from $|H|=1$ to $|H|=170$, by orderly including the nodes with largest degree, one-by-one, to the lesioned set $H$.